OpenStudy (anonymous):

how do i do this ? 4(x + 6) = 6(x − 2)

OpenStudy (anonymous):

expand the brackets first

OpenStudy (anonymous):

all that i know is that it would become 4x+24

OpenStudy (anonymous):

U have to start by expanding the brackets first ...... then u will have 4X+24=6X-12 THEN WILL HAVE 2X=36 X=18

OpenStudy (anonymous):

yep that is correct, what about for the second bracket on the right hand side?

OpenStudy (anonymous):

@kristopher1014 You should expand the parenthesis on both sides 4(x+6) = 4x+24 6(x-2) = 6x-12 4x+24=6x-12 -4x+12 -4x+12 36=2x 18=x

OpenStudy (anonymous):

6x-12

OpenStudy (anonymous):

so its 4x+24=6x+12 but how did u get 12 out of 24 and 4 outta 6?

OpenStudy (anonymous):

6(x-2) = 6x-12

OpenStudy (anonymous):

ok out of 4x+24 i got 6 and out of 6x+12 i got 2 wtf am i doing wrong ?

OpenStudy (anonymous):

4(x+6) = 4(x)+4(6) 6(x-2) = 6(x)=6(-2)

OpenStudy (anonymous):

ok ill just put the answer in, i dont understand where u got that extra 4 but thnks for the help